OpenStudy (anonymous):

What is the area of a circle that has a diameter of 70 cm? Use 22/7 for p.

OpenStudy (anonymous):

22/7*35^2=? can u get it from here?

OpenStudy (anonymous):

no that aint it

OpenStudy (anonymous):

nvrmind it is

OpenStudy (anonymous):

3850cmsquared

OpenStudy (anonymous):

a=pi*r^2 r=1/2*diameter so a= 22/7 * 35^2

OpenStudy (anonymous):

Sid got it